OpenStudy (anonymous):

The graph of the following system of equations is −2x + y = 3 −4x + 2y = 6

OpenStudy (anonymous):

The graph of the following system of equations is −2x + y = 3 −4x + 2y = 6 Overlapping lines Parallel lines Intersecting lines